OpenStudy (anonymous):

Jay Corporation has earned $175,900 after tax. The accountant calculated the return on equity as 12.5 percent. Jay Corporation's stockholders' equity to nearest dollar is:

ganeshie8 (ganeshie8):

i think you need to compute 12.5 % of earnings after tax use the formula x % of something = \[\frac{x}{100}*something\]

ganeshie8 (ganeshie8):

plug in x = 12.5 something = 175,900

ganeshie8 (ganeshie8):

can you do it now ?

OpenStudy (anonymous):

I am trying it now.

ganeshie8 (ganeshie8):

ok.. good :)

OpenStudy (anonymous):

I don't get it?

ganeshie8 (ganeshie8):

x% of something =\[ \frac{x}{100}*something \] 12.5 % of $175,900 = \[ \frac{12.5}{100}*175,900 \] = ?

OpenStudy (anonymous):

21,987.50

OpenStudy (anonymous):

Is that right? and if it were rounded to the nearest dollar it would be $21,988. ??

ganeshie8 (ganeshie8):

wait.. im checking

OpenStudy (anonymous):

ok

ganeshie8 (ganeshie8):

i got the same answer as u :) also u r correct after rounding to nearest dollar it is $21,988 Enjoy..

OpenStudy (anonymous):

ok thats good these were the choices I was givin. $140,720 $14,720 $1,407,200 $140,720,000 None of these

ganeshie8 (ganeshie8):

none of these is ur answer :)

OpenStudy (anonymous):

Thats what I was thinking. But what dose the 100 represent??

ganeshie8 (ganeshie8):

whr ?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

What dose whr mean?

ganeshie8 (ganeshie8):

short cuts.. lol.. sorry. where is 100. . which 100 you asking about ?

OpenStudy (anonymous):

12.5 ---- 100 The 12.5 over 100

ganeshie8 (ganeshie8):

ok.. that 100 represents this 12.5 % in decimal form : 12.5 % = 12.5 over 100 just remove the % and divide it with 100

ganeshie8 (ganeshie8):

very good question btw :)

OpenStudy (anonymous):

OK I got it now! Thank You for your help!! I am sure I will be back shprtly with more questions???? LOL

ganeshie8 (ganeshie8):

yw.. :) :)

ganeshie8 (ganeshie8):

actually, in the above problem $175,900 is 12.5 % of stock holder's equity. problem asks us to compute what is the total stock holder's equity. so the formula becomes : total equity = (175,900/12.5) * 100 = 1407200

OpenStudy (anonymous):

I see. Well I got 8 out of 10 right anyway. :-)

ganeshie8 (ganeshie8):

good to hear :) inspite of this messup... great work :):)

OpenStudy (anonymous):

Thank You!